OpenStudy (abbles):

How is the range of y = 3^x - 4 ..?

OpenStudy (shawn):

(-4, inf)

OpenStudy (abbles):

How is the range of the previously stated equation (-4, infinity)? Shouldn't it be (-3, infinity)? Since the graph of y = 3^x is moved down 4 units and y = 3^x would equal 1 when x = 0 (since all numbers to the 0th power equal 1). 1 minus 4 equals -3... not -4

OpenStudy (abbles):

Right, thanks :) I'm mostly wondering WHY/HOW that is...

OpenStudy (shawn):

well , what is the range of 3^x?

OpenStudy (abbles):

(1, inf) right?

OpenStudy (shawn):

no, it's (0,inf)

OpenStudy (abbles):

3^0 = 1

OpenStudy (abbles):

How is that?

OpenStudy (shawn):

if x approaches negative infinity, y approaches 0+

OpenStudy (shawn):

so numbers like 0.000001 is in the range, yes?

OpenStudy (abbles):

How could the value of y ever equal less than 1 if x^0 = 1... and the exponents can't be negative

OpenStudy (abbles):

I was under the impression the exponents couldn't be negative for some reason... maybe not?

OpenStudy (shawn):

yes, exponent CAN be negative. for example, 3^(-2) = 1/3^2 = 1/9

OpenStudy (shawn):

so do you see why the range of 3^x is (0,inf)? and since 3^x - 4 is just 4 units down, the range is (-4,inf)

OpenStudy (abbles):

Ohhh.. gotcha. So the domain of a^x is (-inf, inf) and the range is (0, inf)?

OpenStudy (abbles):

Why is the range always positive?

OpenStudy (shawn):

because the values of "a" are defined to be 0 < a < 1, or a > 1.

OpenStudy (shawn):

i.e is is positive. A positive number raised to any power is always positive

OpenStudy (abbles):

Oh okay, so the base cannot be negative?

OpenStudy (abbles):

Got it! Thanks for clearing that up

OpenStudy (shawn):

well, base can certainly be negative but there are many problems for example (-3)^(1/3) is defined, but (-3)^(1/2) isn't.

OpenStudy (abbles):

Hmmm

OpenStudy (abbles):

But it's still always positive

OpenStudy (shawn):

the problem is if the base is negative, the function is no longer continuous . What what exponent functions to be continuous

OpenStudy (shawn):

no (-3)^(1/3) isn't positive

OpenStudy (shawn):

anyways, it has something to do with continuity. But that's higher math. For now, just now that the base is always positive, except for 1

OpenStudy (abbles):

Thanks a bunch :) By "higher math", what do you mean? Calculus?

OpenStudy (shawn):

yes, calculus or higher